OpenStudy (helpblahblahblah):

Medal+Fan plz help :) What is the rational number equivalent to 3.12 with a bar over 12? A. 3 and 4 over 33 B. 3 and 8 over 33 C. 3 and 10 over 39 D. 3 and 5 over 39

OpenStudy (helpblahblahblah):

@Opcode can u help?

OpenStudy (opcode):

Do you know what a rational number is?

OpenStudy (helpblahblahblah):

kinda

OpenStudy (opcode):

Okay well do you know what the bar over 12 means?

OpenStudy (helpblahblahblah):

repeating decimal?

hartnn (hartnn):

so where r u stuck! ?

OpenStudy (helpblahblahblah):

on how to make it a fraction

OpenStudy (opcode):

You are correct, the bar over the 12 indicates that. 3.12 has 121212121212 repeating after the decimal point forever.

OpenStudy (opcode):

Do you know how to turn mixed fractions into improper fractions?

OpenStudy (helpblahblahblah):

yes

OpenStudy (helpblahblahblah):

a little

OpenStudy (opcode):

Can you convert: \[3 \dfrac{4}{33}\] Into a improper fraction?

OpenStudy (helpblahblahblah):

\[\frac{ 103 }{ 33 }\] ?

OpenStudy (opcode):

Correct, now to the division that fraction entails. So: \[103 \div 33\]

OpenStudy (helpblahblahblah):

3.12121212...

OpenStudy (opcode):

So the answer choice is?

OpenStudy (helpblahblahblah):

A

OpenStudy (opcode):

Correct!

OpenStudy (helpblahblahblah):

thxx! :D :D

OpenStudy (opcode):

You are welcome. :-)

hartnn (hartnn):

rule of thumb .a bar = a/9 .aa bar = aa/99 .aaa bar = aaa/999 so here .12 bar = 12/99

hartnn (hartnn):

12/99 is 4/33 3.12 bar is 3 4/33 :)

OpenStudy (opcode):

Woah, I did not know about that. Amazing trick @hartnn

OpenStudy (helpblahblahblah):

thx aswell @hartnn :)

hartnn (hartnn):

^_^